# AI Safety in the Eyes of the Downstream Developer: A First Look at Concerns, Practices, and Challenges

### v1

@@ 3.2. Method 2: Survey
Instrument Design. After gathering qualitative results for the RQs, we conducted a survey to triangulate our findings with quantitative insights. We designed our survey to investigate the perceived importance of checking AI safety concerns (RQ1), frequency of adopted actions (RQ2), and the agreement on the challenges (RQ3) at a subtheme granularity. Similar to the design in other studies (Aniche et al., 2018), respondents indicated their agreement and experience with statements derived from the interview analysis subthemes. Given that developers’ practices depend on the concern they are trying to address, for RQ2, we asked about the frequency of adopting certain actions separately for each of the three main themes from RQ1. Due to the large number of practice and AI safety concern combinations, we selected practices mentioned by at least half of the participants to minimise fatigue and ensure response quality. Detailed scales are shown in Table 1. We allowed users to respond “I do not know” or “Does not apply” for all questions. The survey included 43 statements and took approximately 15 minutes to complete.
@@
Recruitment. We released our survey on the Prolific platform. To filter participants without prior experience in developing AI-based software, we applied a pre-defined filter from Prolific, requiring participants to be employed in the Information Technology (IT) sector, as in a previous study (Gao et al., 2023). We excluded participants who reported no experience in deep learning, software engineering, or in developing AI-based software. Following Schmidt et al., we asked participants to describe one of their last three AI-based software projects and their roles (Schmidt et al., 2023). Two authors independently reviewed the authenticity of the project and job role description to ensure a high quality of recruited participant (Schmidt et al., 2023). Of the 128 participants recruited, we screened out 27 based on experience questions, and 15 more based on project descriptions, resulting in 86 valid participants. For demographics, 51% had built 1-5 AI-based software projects using PTMs, 26% built 5-10, 13% built 10-25, and 12% built more than 25. For deep learning experience, 47% had 1-2 years, 38% had 3-5 years, 9% had 6-10 years, and 6% had more than 10 years. For software engineering experience, 40% had 1-2 years, 21% had 3-5 years, 21% had 6-10 years, and 20% had more than 10 years.
@@
Analysis. Similar to previous studies (Gao et al., 2023), we converted the responses to a scale of 1 to 4 and applied the Wilcoxon signed-rank test (Rey and Neuhäuser, 2011) with Cohen’s d effect size (Becker, 2000) for pairs of practices under different AI safety concerns to identify whether developers adopt strategies differently when addressing different AI safety concerns. This test is well-suited to our data, given the ordinal nature of the scores and the dependency between paired scores, as they were provided by the same group of participants. In cases of missing data, we included only pairs where both responses were available.
@@ 3.3. Method 3: AI Incident Analysis
While interviews and surveys capture downstream developers’ perspectives on AI safety, they do not reveal whether these concerns align with real-world failures. To address this gap, we analysed incidents from the AI Incident Database. By triangulating these perspectives, we identify potential mismatches between developer focus and real-world safety outcomes.
Data Source. We downloaded all 874 incidents from the AI Incident Database, which indexes a collective history of harms or near-harms caused by the deployment of artificial intelligence systems in the real world (inc, [n. d.]). Managed by a board of responsible AI experts, the database compiles public submissions and is recognised as a valuable resource for identifying AI system failures (Anandayuvaraj et al., 2024). Each entry in the database includes a description of the incident.
@@ 4. Results
Here we cover our results, organised by RQ. Table 3 summarises.
Table 3. Summary of our results: sub-themes and frequency by interview participants, survey respondent, and AI incident. The bolded phrases of the sub-themes are used as labels in the figures for survey results (Figures 3, 4 and 5).
@@ 4.1. RQ1: AI Safety Concerns
Developers’ AI safety concerns include three themes, namely (1) data concerns, (2) model behaviour concerns, and (3) miseuse and exploitation concerns. Figure 3 illustrates the RQ1 survey results.
Figure 3. Survey results on importance of AI safety concerns.
@@ 4.2. RQ2: Practices when approaching AI safety
Figure 4. Survey results on frequency of AI safety practices. (* indicates statistical significance vs. other concerns)
@@
Theme 2: Implementing technical safety guards involves adopting methods to mitigate AI safety issues during development. Improve safety assurance by iteratively refining models and systems requires developers to adjust models, evaluate safety factors, and iterate until satisfactory results are achieved. The fine-tuning process may be examined, as P13 mentioned: “first we tried to see if we did something wrong in the fine-tuning process”. This also includes examining fine-tuning processes, switching models or checkpoints, and retraining on improved datasets. For example, P17 fine-tuned with a more diverse accent dataset to enhance speech recognition performance for rural accents. For LLMs, practitioners also refine prompts to prevent undesired content generation. P13 reflected: “in terms of politeness that I mentioned, sometimes we try to change the prompt we are using for the model to prevent it from talking things we don’t want to do”. Our survey revealed significant differences in responses depending on the AI safety context. Developers performed iterative model and system improvement significantly more frequently when addressing model behaviour risks (3.24) compared to data (3.00) and misuse and exploitation (2.98) concerns, with p-values of 0.017 and 0.004, and effect sizes of 0.32 and 0.33, respectively. These results suggest that different AI safety concerns have a medium effect on developers’ frequency of adopting iterative model and system improvement.
@@
The previous efforts relies on the actions to ensure dataset quality. One approach is to incorporate higher-quality data into fine-tuning dataset for the prioritised AI safety concerns. For example, since P17’s speech recognition product targets children in multiple rural areas, they partnered with local organisations to gather diverse accents, enriching their fine-tuning data for model improvement. Filtration methods are also used to remove biased, violent, or personally sensitive content in fine-tuning datasets. For example, when developing AI products for banks, P7 masked customers’ sensitive information before feeding the data into the model, as data privacy is a priority in the financial industry (Akash et al., 2024). These filtration methods involve manual annotation by human reviewers or automated filtering. Also, sampling and data engineering techniques are employed to enhance datasets to provide better inputs aligned with the prioritised AI safety concerns. Our survey shows this practice had the highest average score of 3.25, aligning with the high importance placed on data concerns.
@@
Filtering model input and output acts as an external safeguard against AI safety violations, using both manual rules and automated filters. For example, P15 applied filters to detect inappropriate content, while P12 used regular expressions to remove privacy-sensitive information from model-generated text. For LLMs, P11 highlighted using Llama Guard (Inan et al., 2023) as a post-processing step for model output. Following security practices in the development of AI systems relates to particular processes needed in the development. Several participants highlighted local solutions, such as on-premise data storage and training pipelines to prevent data leakage. For example, P12 addressed data storage concerns by stating: “The data leakage can happen and to resolve it, … on premise the installation of databases”. Other security measures include encrypting data and models, maintaining backups, and selectively publishing checkpoint versions. Based on our survey data, this subtheme is adopted frequently, receiving an average score of 3.24.
@@
Theme 3: Maintaining documentation and expert consultation contains several procedural efforts. Maintain clear documentation refers to the process of providing documentation for both internal review and downstream users to ensure transparency. P9 mentioned that all processes, including decisions on PTM selection, model robustness, and efforts to mitigate risks, need to be clearly documented for company’s audit. Participants also made efforts in recording model evaluations for various AI safety concerns, providing guidelines for responsible interaction with the software, and acknowledging potential AI safety limitations in the documentation. Additionally, documentation on disclaimers was mentioned, with P12 highlighting its use as a way to bypass responsibility for AI safety breaches. Our survey respondents reported varying and slightly lower frequencies of documentation practices across different concerns: data (2.79), model behaviour (3.04), and misuse and exploitation (2.95). Documentation was significantly more frequent for model behaviour concerns than for data concerns (p = 0.017, effect size = 0.30), suggesting a medium impact of concern type on documentation frequency. Overall, documentation ranked among the least frequently adopted AI safety practices.
@@
Meanwhile, cross-functional collaboration for AI safety assurance involves the cooperative efforts to address AI safety concerns, including discussions on risks, collaboration among multiple teams or organisations to collect datasets, test AI models for safety concerns, deploy software securely, and maintenance over time. Moreover, given the expertise gap, AI software developers collaborate with domain experts throughout the development process to analyse risks, build evaluation ground truth datasets, and conduct further evaluations. For example, when handling sensitive customer data, P13 reflected: “I’m not an expert in that field. I just talk with the legal team or the compliance team and if they say no, I just stop … we are working constantly with that team”. This subtheme received an average score of 3.09, indicating practitioners’ broad adoption.
@@
Theme 4: Regular monitoring and assessment for safety concerns involves evaluating AI safety before release and continuous monitoring systems after deployment. A key step is to develop evaluation dataset for AI model safety, with diverse efforts observed, from simply using open-source data to crowdsourcing or even expert consultations. One way to construct evaluation dataset is to reuse existing evaluation dataset on open-source platforms, as P4 mentioned collecting testing documents related to sensitive topics on GitHub. Efforts are also spent refining existing datasets by reformulating certain details. For instance, P6 described adapting culturally American-centric question-answer pairs to reflect the traditions of other culture to evaluate models. When assessing model’s reliability, robustness, or data leakage, practitioners often use ground-truth or production data. For hallucination evaluation, P16 reflected: “the customer support case, …, we will have transcripts of customer support conversations”. Last, P13 also exploit the possibility of using LLM to generate test cases to evaluate model AI safety: “we have used some generative AI models like GPT to maybe help us generate some questions”. Our survey results showed this practice received an average score of 3.01 for its frequency of adoption, indicating it is widely used for AI safety assessments.
@@
After that, developers would test and evaluate AI model safety. Diverse methods were observed among the participants. Some efforts are preliminary, involving only probing the model with small samples and manual inspection,, such as P4:“we randomly select some questions and test the LLM to see if it response with safe answers”. Automatic evaluation using traditional metrics such as Rouge (Lin, 2004) was also mentioned. In particular, LLM-as-judge (Zheng et al., 2023) is used by developers to evaluate model safety factors such as hallucinations, bias, and harmful generation, as reflected by P16. Also, a hybrid approach combines manual and automated effort, as P18 described for privacy evaluation: “then flagging low scoring examples in a development set for … human evaluation”. Our survey indicates that developers conduct model safety testing and evaluation significantly more frequently when addressing model behaviour risks (3.22) compared to data concerns (2.95) and misuse and exploitation concerns (2.94), with p-values of 0.007 and 0.004 and effect sizes of 0.40 and 0.41, respectively. These results suggest that different AI safety concerns have a moderate effect on the frequency with which developers test and evaluate AI model safety.
@@
Highlight: Practitioners’ approaches to AI safety vary in comprehensiveness. Documentation and feedback collection are less frequently adopted, while fewer interview participants consider understanding safety concerns as an initial step.

### v3

@@ 3.2. Method 2: Survey
Instrument Design. After gathering qualitative results for the RQs, we conducted a survey to triangulate our findings with quantitative insights. We designed our survey to investigate the perceived importance of checking AI safety concerns (RQ1), frequency of adopted actions (RQ2), and the agreement on the challenges (RQ3) at a subtheme granularity. We conducted five pilot studies internally with people with experience in the development of AI-based software. Based on their feedback, we refined the clarity of the questions and adjusted the survey length. Similar to the design in other studies (Aniche et al., 2018), respondents indicated their agreement and experience with questions derived from the interview analysis subthemes. Given that developers’ practices depend on the concern they are trying to address, for RQ2, we asked about the frequency of adopting certain actions separately for each of the three main themes from RQ1. Due to the large number of practice and AI safety concern combinations, we selected practices mentioned by at least half of the participants to minimise fatigue and ensure response quality. Detailed scales are shown in Table 1. We allowed users to respond “I do not know” or “Does not apply” for all questions. The survey included 43 questions and took approximately 15 minutes to complete.
@@
Recruitment. We released our survey on the Prolific platform. To filter participants without prior experience in developing AI-based software, we applied a pre-defined filter from Prolific, requiring participants to be employed in the Information Technology (IT) sector, as in a previous study (Gao et al., 2023). We excluded participants who reported no experience in deep learning, software engineering, or in developing AI-based software. Following Schmidt et al., we asked participants to describe one of their last three AI-based software projects and their roles (Schmidt et al., 2023). Two authors independently reviewed the authenticity of the project and job role description to ensure a high quality of recruited participant (Schmidt et al., 2023).
Data Validity Checks. Of the 130 participants recruited, we screened out 27 based on experience questions, and 15 more based on project descriptions, resulting in 88 valid participants. We also included two attention check questions to ensure data quality (Beach, 1989), which required participants to select certain choices in the survey. For example, one of the attention questions instructed the participants to select the choice of “neutral” as the response. Participants who failed either of the attention checks were considered invalid. This further excluded two responses, resulting in a final total of 86 valid survey submissions.
@@
Analysis. Similar to previous studies (Gao et al., 2023), we converted the responses to a scale of 1 to 4 for RQ2, and a scale of 1 to 5 for RQ1 and RQ3 to measure the frequency of adoption for practices, and perceived importance of each AI safety concern and challenge, respectively. To identify whether developers adopt strategies differently when addressing different AI safety concerns, we applied the Wilcoxon signed-rank test (Rey and Neuhäuser, 2011) with Bonferroni correction (Weisstein, 2004) and Cohen’s d effect size (Becker, 2000) for pairs of practices under different AI safety concerns. This test is well-suited to our data, given the ordinal nature of the scores and the dependency between paired scores, as they were provided by the same group of participants. In cases of missing data, we included only pairs where both responses were available. For demographics, 51% had built 1-5 AI-based software projects using PTMs, 26% built 5-10, 13% built 10-25, and 12% built more than 25. For deep learning experience, 47% had 1-2 years, 38% had 3-5 years, 9% had 6-10 years, and 6% had more than 10 years. For software engineering experience, 40% had 1-2 years, 21% had 3-5 years, 21% had 6-10 years, and 20% had more than 10 years.
@@ 3.3. Method 3: AI Incident Analysis
While interviews and surveys capture downstream developers’ perspectives on AI safety, they do not reveal whether these concerns align with real-world failures. To address this gap, we analysed incidents from the AI Incident Database (AIID) (McGregor, 2021). By triangulating these perspectives, we identify potential mismatches between developer focus and real-world safety outcomes.
Data Source. Managed by a board of responsible AI experts, the AI incident database compiles public submissions and provides a collective memory of failings in intelligent systems (McGregor, 2021). The database indexes a collective history of harms or near-harms caused by the deployment of artificial intelligence systems in the real world (inc, [n. d.]). Each entry includes a description of the incident and multiple reports linked to related news coverage. The reports are sampled to ensure comprehensive coverage of each incident and to enhance the discoverability of incidents by users. We downloaded all 874 incidents from the AI Incident Database in 13 Jan 2025.
@@
3.4. Characteristics of Three Data Sources
In this subsection, we describe how we used the three data sources to answer our research questions.
The interpretation of the AI safety definition is shared across the three data sources. We started with a qualitative analysis of the interview transcripts to derive subthemes for each RQ. After that, we formulated the subthemes that emerged from our interview data analysis into questions (e.g., How important is it to mitigate risks of the software being used irresponsibly by end-users?). Therefore, we did not need to provide a high-level AI safety definition to our survey participants, but only specific questions that were derived from the analysis of the interviews. Similarly, we used the subthemes derived from the interviews to annotate the incidents in the AI incident database to identify which specific AI safety concerns each incident exemplifies.
To assess the relevance and coverage of the three data sources, we compared the demographics of participants recruited for the interview and survey. We also compared the application domain and safety criticality of the applications that our interview participants and survey participants worked on, along with the applications mentioned in the AI incident database. Regarding the demographics of the recruited participants, interview participants had an average of 5.6 years of SE experience and 4.1 years of DL experience, with a standard deviation of 5.9 and 2.7, respectively. For the survey participants, we collected their years of experience in the form of intervals. Therefore, we take the medium value of each interval for representation to facilitate comparison. For the survey participants, they have on average 5.6 years of SE experience and 3.6 years of DL experience, with a standard deviation of 3.1 and 2.5, respectively. In other words, the experience in SE are comparable between two groups, while interview participants have slightly more years of experience than survey participants in DL.
Regarding the application domain and safety criticality of the developed applications, Table 3 summarises the application domains referenced by both interview and survey participants, along with the criticality of the systems they described. The table also includes the corresponding information from the records in the AI incident database. As some participants described software combining multiple components, such as computer vision and natural language processing, and some survey responses or AI incident descriptions did not clearly specify which models were reused, the totals for CV and NLP applications may not sum to 100%. Each mentioned system was annotated as either safety-critical or non-safety-critical, following the definition by Knight (Knight, 2002), where a safety-critical system is one whose failure could result in loss of life, significant property damage, or environmental harm. During annotation, we adhered to this guideline by focusing on the system’s originally intended purpose rather than the specific context of use. For example, failures in autonomous vehicle systems are considered safety-critical because their intended purpose inherently involves risks to people. In contrast, AI game agents’ inappropriate behaviours cause no direct harm to individuals, and are thus not classified as safety-critical.
Table 3 shows that NLP-based applications are discussed by over 70% of the participants in both the interview and survey, while CV-based applications are discussed less. The percentage of safety-critical systems fluctuated around 10% for the two data sources. In comparison, there were more CV-based applications mentioned in the AI incident database, and safety-critical systems were also more common there, at around 20%.
In summary, the three data sources share a consistent interpretation of AI safety concepts, and we did not specifically target safety-critical systems in our study, as reflected by the low percentages across all three data sources. Although the participants’ expertise levels varied between interviews and survey, the applications they discussed were similar in terms of domain and safety criticality. In the AI incident database analysis, the application domains show more variability; however, the majority of the software applications were also not safety-critical.
Table 3. Summarisation of Application and Criticality of Three Data Sources
Data Source CV application NLP application Safety Critical System
Interview 37.5% 81.3% 12.5%
Survey 15.1% 70.9% 9.3%
AI-incident database 43.0% 38.1% 20.0%
In section 4 we provide the results to answer each of the RQs. For each RQ, we first present the qualitative analysis results from the interview participants and then quantitative findings from the survey participants. As the AI incident database only reveals the real-world failures mapping to the AI safety concerns, it is discussed in section 4.1 as a dimension for RQ1. In section 5, we structure the discussion around different stakeholder groups. The qualitative results from interviews form the foundation, while survey scores provide insight into how downstream developers perceive the importance of various concerns, the frequency with which they adopt practices, and the severity of challenges. Additionally, our analysis of the AI Incident Database serves as a proxy to assess whether these concerns are being adequately addressed in practice.
@@ 4. Results
Here we cover our results, organised by RQ. Table 4 summarises the interview, survey and AI incident database results.
Table 4. Summary of our results: sub-themes and frequency by interview participants, survey respondent, and AI incident. The bolded phrases of the sub-themes are used as labels in the figures for survey results (Figures 4, 5 and 6). The survey score column represents the average of the survey scores, while the detailed distribution is shown in Figures 4, 5 and 6.
@@ 4.1. RQ1: AI Safety Concerns
Developers’ AI safety concerns include three themes, namely (1) data concerns, (2) model behaviour concerns, and (3) misuse and exploitation concerns. Figure 4 illustrates the RQ1 survey results.
Figure 4. Survey results on importance of AI safety concerns. The full details about the distribution is in our replication package. section 8
@@ 4.2. RQ2: Practices when approaching AI safety
Figure 5. Survey results on frequency of AI safety practices. The full details about the distribution is in our replication package section 8
@@
Theme 2: Implementing technical safety guards involves adopting methods to mitigate AI safety issues during development. Improve safety assurance by iteratively refining models and systems requires developers to adjust models, evaluate safety factors, and iterate until satisfactory results are achieved. The fine-tuning process may be examined, as P13 mentioned: “first we tried to see if we did something wrong in the fine-tuning process”. This also includes examining fine-tuning processes, switching models or checkpoints, and retraining on improved datasets. For example, P17 fine-tuned with a more diverse accent dataset to enhance speech recognition performance for rural accents. For LLMs, practitioners also refine prompts to prevent undesired content generation. P13 reflected: “in terms of politeness that I mentioned, sometimes we try to change the prompt we are using for the model to prevent it from talking things we don’t want to do”. Our survey revealed significant differences in responses depending on the AI safety context. Developers performed iterative model and system improvement significantly more frequently when addressing model behaviour risks (3.24) compared to misuse and exploitation (2.98) concerns, with p-values of and effect sizes of 0.33. These results suggest that different AI safety concerns have a medium effect on developers’ frequency of adopting iterative model and system improvement. Only 13% of the participants put low priority on this practice when dealing with model behaviour concerns, which is lower compared to the scenarios when dealing with data or misuse and exploitation concerns.
@@
The previous efforts relies on the actions to ensure dataset quality. One approach is to incorporate higher-quality data into fine-tuning dataset for the prioritised AI safety concerns. For example, since P17’s speech recognition product targets children in multiple rural areas, they partnered with local organisations to gather diverse accents, enriching their fine-tuning data for model improvement. Filtration methods are also used to remove biased, violent, or personally sensitive content in fine-tuning datasets. For example, when developing AI products for banks, P7 masked customers’ sensitive information before feeding the data into the model, as data privacy is a priority in the financial industry (Akash et al., 2024). These filtration methods involve manual annotation by human reviewers or automated filtering. Also, sampling and data engineering techniques are employed to enhance datasets to provide better inputs aligned with the prioritised AI safety concerns. Our survey shows this practice had the highest average score of 3.25, aligning with the high importance placed on data concerns. More than 80% of the survey participants consistently adopt this strategy as a regular or priority practice when handling all the mentioned AI safety concerns.
@@
Filtering model input and output acts as an external safeguard against AI safety violations, using both manual rules and automated filters. For example, P15 applied filters to detect inappropriate content, while P12 used regular expressions to remove privacy-sensitive information from model-generated text. For LLMs, P11 highlighted using Llama Guard (Inan et al., 2023) as a post-processing step for model output. Following security practices in the development of AI systems relates to particular processes needed in the development. Several participants highlighted local solutions, such as on-premise data storage and training pipelines to prevent data leakage. For example, P12 addressed data storage concerns by stating: “The data leakage can happen and to resolve it, … on premise the installation of databases”. Other security measures include encrypting data and models, maintaining backups, and selectively publishing checkpoint versions. Based on our survey data, this subtheme is adopted frequently, receiving an average score of 3.24, and it was considered a regular to priority practice by more than 80% of the participants across all AI safety concern contexts.
@@
Theme 3: Maintaining documentation and expert consultation contains several procedural efforts. Maintain clear documentation refers to the process of providing documentation for both internal review and downstream users to ensure transparency. P9 mentioned that all processes, including decisions on PTM selection, model robustness, and efforts to mitigate risks, need to be clearly documented for company’s audit. Participants also made efforts in recording model evaluations for various AI safety concerns, providing guidelines for responsible interaction with the software, and acknowledging potential AI safety limitations in the documentation. Additionally, documentation on disclaimers was mentioned, with P12 highlighting its use as a way to bypass responsibility for AI safety breaches. Our survey respondents reported varying and slightly lower frequencies of documentation practices across different concerns: data (2.79), model behaviour (3.04), and misuse and exploitation (2.95). Overall, documentation ranked among the least frequently adopted AI safety practices, with around 24% to 34% of the participants de-prioritising this task. This indicates a relatively lower focus on documentation efforts from the downstream developers’ side.
@@
Meanwhile, cross-functional collaboration for AI safety assurance involves the cooperative efforts to address AI safety concerns, including discussions on risks, collaboration among multiple teams or organisations to collect datasets, test AI models for safety concerns, deploy software securely, and maintenance over time. Moreover, given the expertise gap, AI software developers collaborate with domain experts throughout the development process to analyse risks, build evaluation ground truth datasets, and conduct further evaluations. For example, when handling sensitive customer data, P13 reflected: “I’m not an expert in that field. I just talk with the legal team or the compliance team and if they say no, I just stop … we are working constantly with that team”. This subtheme received an average score of 3.09, with fewer than 30% of participants indicating it was a low priority, suggesting broad adoption among practitioners.
@@
Theme 4: Regular monitoring and assessment for safety concerns involves evaluating AI safety before release and continuous monitoring systems after deployment. A key step is to develop evaluation dataset for AI model safety, with diverse efforts observed, from simply using open-source data to crowdsourcing or even expert consultations. One way to construct evaluation dataset is to reuse existing evaluation dataset on open-source platforms, as P4 mentioned collecting testing documents related to sensitive topics on GitHub. Efforts are also spent refining existing datasets by reformulating certain details. For instance, P6 described adapting culturally American-centric question-answer pairs to reflect the traditions of other culture to evaluate models. When assessing model’s reliability, robustness, or data leakage, practitioners often use ground-truth or production data. For hallucination evaluation, P16 reflected: “the customer support case, …, we will have transcripts of customer support conversations”. Last, P13 also exploit the possibility of using LLM to generate test cases to evaluate model AI safety: “we have used some generative AI models like GPT to maybe help us generate some questions”. Our survey results showed this practice received an average score of 3.01 for its frequency of adoption, with 73% to 76% of participants reporting it as a regular or prioritized activity, indicating it is widely used for AI safety assessments.
@@
After that, developers would test and evaluate AI model safety. Diverse methods were observed among the participants. Some efforts are preliminary, involving only probing the model with small samples and manual inspection,, such as P4:“we randomly select some questions and test the LLM to see if it response with safe answers”. Automatic evaluation using traditional metrics such as Rouge (Lin, 2004) was also mentioned. In particular, LLM-as-judge (Zheng et al., 2023) is used by developers to evaluate model safety factors such as hallucinations, bias, and harmful generation, as reflected by P16. Also, a hybrid approach combines manual and automated effort, as P18 described for privacy evaluation: “then flagging low scoring examples in a development set for … human evaluation”. Our survey indicates that developers conduct model safety testing and evaluation significantly more frequently when addressing model behaviour risks (3.22) compared to data concerns (2.95) and misuse and exploitation concerns (2.94), with p-values of 0.007 and 0.004 and effect sizes of 0.40 and 0.41, respectively. 16% of the participants put low priority on this practice when dealing with model behaviour concerns, while this number increased to 23% and 27% within the context of data concerns and misuse and exploitation concerns, respectively. These results suggest that different AI safety concerns have a moderate effect on the frequency with which developers test and evaluate AI model safety.
@@
Highlight: Practitioners’ approaches to AI safety vary in comprehensiveness. Documentation and feedback collection are less frequently adopted, and few interview participants consider understanding safety concerns as an initial step.